# suppSupplementary References

### v2

@@ Acknowledgements
If the values of the generalized quantile functions $F^{-1}$ and $G^{-1}$ are known at levels $0\leq\alpha_{1}<\dots<\alpha_{K}\leq 1$ and $0\leq\beta_{1}<\beta_{2}<\ldots<\beta_{L}\leq 1$, respectively, the Cramér distance can be approximated as
$\operatorname{CD}(F,G)\approx 2\sum_{i=1}^{K}\sum_{j=1}^{L}\frac{1+\mathbbm{1}\{\alpha_{i}\neq\beta_{j}\}}{2}\frac{\alpha_{i+1}-\alpha_{i-1}}{2}\frac{\beta_{j+1}-\beta_{j-1}}{2}\chi\left(\alpha_{i},\beta_{j}\right)\left|F^{-1}(\alpha_{i})-G^{-1}\left(\beta_{j}\right)\right|,$
where $\alpha_{0}=\beta_{0}=0$ and $\alpha_{K+1}=\beta_{L+1}=1$. If the quantiles bound central intervals (i.e., $\alpha_{i}=1-\alpha_{K+1-i}$ and $\beta_{j}=1-\beta_{L+1-j}$ holds for $i=1,\dots,K$ and $j=1,\dots,L$), then the components can be approximated similarly as
$\operatorname{Shift}_{+}^{\operatorname{CD}}(F,G)\approx\sum_{i=1}^{\lceil K/2\rceil}\sum_{j=1}^{\lceil L/2\rceil}(1+\mathbbm{1}\{\alpha_{i}\neq 0.5,\beta_{j}\neq 0.5\})\frac{\alpha_{i+1}-\alpha_{i-1}}{2}\frac{\beta_{j+1}-\beta_{j-1}}{2}\\ \cdot\left(\Big{[}\min\Big{\{}F^{-1}\left(\alpha_{K+1-i}\right)-G^{-1}\left(\beta_{L+1-j}\right),F^{-1}\left(\alpha_{i}\right)-G^{-1}\left(\beta_{j}\right)\Big{\}}\Big{]}_{+}+\left[F^{-1}\left(\alpha_{i}\right)-G^{-1}\left(\beta_{L+1-j}\right)\right]_{+}\right)$
and
$\operatorname{Disp}_{+}^{\operatorname{CD}}(F,G)\approx\sum_{i=1}^{\lceil K/2\rceil}\sum_{j=1}^{\lceil L/2\rceil}(1+\mathbbm{1}\{\alpha_{i}\neq 0.5,\beta_{j}\neq 0.5\})\frac{1+\mathbbm{1}\{\alpha_{i}\neq\beta_{j}\}}{2}\frac{\alpha_{i+1}-\alpha_{i-1}}{2}\frac{\beta_{j+1}-\beta_{j-1}}{2}\\ \cdot\Big{[}\Big{(}F^{-1}\left(\alpha_{K+1-i}\right)-G^{-1}\left(\beta_{L+1-j}\right)\Big{)}-\Big{(}F^{-1}\left(\alpha_{i}\right)-G^{-1}\left(\beta_{j}\right)\Big{)}\Big{]}_{+}$
The minus counterparts are approximated using the above formulas via the familiar symmetries $\operatorname{Shift}_{-}^{\operatorname{CD}}(F,G)=\operatorname{Shift}_{+}^{\operatorname{CD}}(G,F)$ and $\operatorname{Disp}_{-}^{\operatorname{CD}}(F,G)=\operatorname{Disp}_{+}^{\operatorname{CD}}(G,F)$. Formulas without the correction factor of $\frac{1}{2}$ that is applied in the case of equal quantile levels $\alpha_{i}=\beta_{j}$ tend to overestimate the CD and the corresponding dispersion components. The supplementary replication code provides an implementation of the approximation formulas, which sometimes served as a good alternative to using numeric integration techniques for the double integrals in (15)–(16).
S4.2 Binned probability distributions
In the application from Section 5.2, the assigned bin probabilities together with the bin boundaries (i.e., the values separating the bins) identify the distribution and quantile functions at these boundary values. However, no further information on the quantile functions can be inferred from the responses.
In Figure 13, we illustrate the points where the distributions are known by the solid points in the quantile spread plots. To compute the Cramér and Wasserstein distances and their associated decompositions in this case, the quantile functions need to be approximated. For this, we adopt a simple linear interpolation between the known quantiles of the distributions as illustrated in Figure 13 with the straight lines connecting the points. In terms of the underlying histogram forecasts, this corresponds to uniformly distributing the probability mass within each bin.
@@
For real numbers $A,B\in\mathbb{R}$ and $[A]_{+}:=\max(A,0)$, we have
@@
  • If $A,B>0$, the statement is immediate as $[A]_{+}+[B]_{+}=A+B$ and $[A+B]_{+}+[\min\{A,-B\}]_{+}+[\min\{-A,B\}]_{+}=A+B+0+0$.
  • If $A,B\leq 0$, then $[A]_{+}+[B]_{+}=0+0$ and $[A+B]_{+}+[\min\{A,-B\}]_{+}+[\min\{-A,B\}]_{+}=0+0+0$.
  • If $A>0$ and $B\leq 0$, we get that $[A]_{+}+[B]_{+}=A+0=A$. If additionally $A\geq-B$, we get $[A+B]_{+}+[\min\{A,-B\}]_{+}+[\min\{-A,B\}]_{+}=(A+B)+(-B)+0=A$. Similarly, if $A<-B$, we get $[A+B]_{+}+[\min\{A,-B\}]_{+}+[\min\{-A,B\}]_{+}=0+A+0=A$.
@@
where the third equality is obtained by a simple change of variables and the last equality is obtained by applying Lemma S5.1 to $[A(\alpha)]_{+}+[-B(\alpha)]_{+}$ and $[-A(\alpha)]_{+}+[B(\alpha)]_{+}$.
@@
$=\int_{-\infty}^{y}\int_{0}^{1}\underbrace{2(-\mathbbm{1}\{x\geq F^{-1}(\tau)\})(\mathbbm{1}\{y\leq F^{-1}(\tau)\}-\tau)}_{\neq 0,\text{ only if}F^{-1}(\tau)\leq x\leq y\text{ for }x\in(-\infty,y]}\,\mathrm{d}\tau\,\mathrm{d}x+\int_{y}^{\infty}\int_{0}^{1}\underbrace{2(\mathbbm{1}\{x<F^{-1}(\tau)\})(\mathbbm{1}\{y\leq F^{-1}(\tau)\}-\tau)}_{\neq 0,\text{ only if }y\leq x<F^{-1}(\tau)\text{ for }x\in[y,\infty)}\,\mathrm{d}\tau\,\mathrm{d}x$
@@
$+$ $\underbrace{\int_{\tau}^{1}(-\tau)(F^{-1}(\tau)-G^{-1}(\tau))\,\mathrm{d}\xi}_{=-(1-\tau)\tau(F^{-1}(\tau)-G^{-1}(\tau))}+\int_{\tau}^{1}\underbrace{\chi(\tau,\xi)(F^{-1}(\tau)-G^{-1}(\xi))}_{\geq 0\text{ as }\xi\geq\tau}\,\mathrm{d}\xi\,\mathrm{d}\tau$
@@
$f\left(\tfrac{1-\alpha}{2},\tfrac{1-\beta}{2}\right)$ =$=\begin{cases}0,&\text{if }\alpha\leq\beta\text{ and }A\geq 0\\ -A,&\text{if }\alpha\leq\beta\text{ and }A\leq 0\\ A,&\text{if }\alpha\geq\beta\text{ and }A\geq 0\\ 0,&\text{if }\alpha\geq\beta\text{ and }A\leq 0\end{cases}$
@@
Notice that for the two minus components with subscript ‘--’, changing the roles of $F$ and $G$ merely changes the sign of $A$ and $B$. ∎
@@
We start by proving Propositions 3.7 and 3.8 as these help to simplify the proofs of some of the basic properties given in Section 3.1. Notice that the proofs of Propositions 3.7 and 3.8 do not require any of the results from Section 3.1 apart from the obvious symmetry given by Proposition 3.1.
@@
    From Proposition 3.8 (shown above), it follows that $\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)=0$ is zero as well for all $p\in\mathbb{N}$.
@@
By Propositions 3.7 and 3.8 (proved above), it suffices to show the equivalences in (a) and (b) for the AVM.
@@
    is zero. Proposition 3.8 implies that $\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)=0$ for all $p\in\mathbb{N}$.
@@
    =$=\begin{cases}\operatorname{AVM}(F_{m_{G}-m_{F}},G),&\text{if }\{\pm=+\text{ and }s_{F}\geq s_{G}\}\text{ or }\{\pm=-\text{ and }s_{F}<s_{G}\},\\ 0,&\text{if }\{\pm=-\text{ and }s_{F}\geq s_{G}\}\text{ or }\{\pm=+\text{ and }s_{F}<s_{G}\}.\end{cases}$
@@
    =$\quad=\begin{cases}\operatorname{AVM}(F,G)-\operatorname{AVM}(F_{m_{G}-m_{F}},G),&\text{if }\{\pm=+\text{ and }m_{F}\geq m_{G}\}\text{ or }\{\pm=-\text{ and }m_{F}<m_{G}\},\\ 0,&\text{if }\{\pm=-\text{ and }m_{F}\geq m_{G}\}\text{ or }\{\pm=+\text{ and }m_{F}<m_{G}\}.\end{cases}$
@@
Let $\widetilde{m}=m_{F}-m_{G}$ be the difference between the central medians of $F$ and $G$ (as defined prior to Proposition 3.3). We assume w.l.o.g. that $\widetilde{m}\geq 0$ (otherwise, we exchange $F$ and $G$). Then, the minus shift components are zero by Proposition 3.3. Furthermore, we define $f(\tau)=F^{-1}(\tau)-G^{-1}(\tau)$ for $\tau\in(0,1)$. By symmetry of $F$ and $G$, the function $f$ is (almost surely) symmetric as well, that is, $f(\tau)=2\widetilde{m}-f(1-\tau)$ holds for almost all $\tau\in(0,1)$. Let $A=\{\tau\mid f(1-\tau)>f(\tau)\geq 0\}\cup\{\tau<\tfrac{1}{2}\mid f(1-\tau)=f(\tau)\geq 0\}$. Then, the plus shift component of the $p$-Wasserstein distance is twice the integral of $f$ over $A$ as a change of variables yields $\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)=\int_{0}^{1}\left[\min\left\{f(\tau)^{[p]},f(1-\tau)^{[p]}\right\}\right]_{+}\,\mathrm{d}\tau=2\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau$. Let $\overline{A}=\{1-\tau\mid\tau\in A\}$, $B=\{\tau\mid f(\tau)\geq 0>f(1-\tau)\}$, and $\overline{B}=\{1-\tau\mid\tau\in B\}$. By the (almost sure) symmetry of $f$, the values $\tau$ such that $0>f(\tau)$ and $0>f(1-\tau)=2\widetilde{m}-f(\tau)$ form a null set in $(0,1)$, and so does the complement of the disjoint union $A\cup\overline{A}\cup B\cup\overline{B}$ in $(0,1)$. Note that the symmetry of $f$ yields the inequalities
$\widetilde{m}$ $\geq f(\tau)$ for almost all $\quad\text{for almost all }\tau\in A$ (S4)
$f(\tau)$ $\geq 2\widetilde{m}$ for almost all $\quad\text{for almost all }\tau\in B$ (S5)
$f(1-\tau)$ $\geq\widetilde{m}$ for almost all $\quad\text{for almost all }\tau\in A$ (S6)
$f(\tau)$ $\geq-f(1-\tau)$ for almost all $\quad\text{for almost all }\tau\in B$ (S7)
We use these four a.s. (almost sure with respect to the Lebesgue measure) inequalities to derive the inequality

### v3

@@ Acknowledgements
for the AVM and its decomposition, and analogously for the CD,
[…]
S4.4 Bounds
In this subsection, we provide bounds for distances and components in settings with incomplete distributions. These bounds are used to study the error of approximations such as the ones proposed above in the next section. For ease of exposition, we only consider a simplified setting. However, generalizations to the more general settings discussed in the previous section are straightforward. As with the approximation of the previous section, the bounds are obtained from certain discrete distributions that are derived from the given quantiles to which the formulas from Section S4.1 are applied. Notably, we only treat the case of distributions with bounded support as otherwise the unknown tail behavior could theoretically inflate the distances arbitrarily. To circumvent this issue when considering distributions with unbounded support, one may consider a censored distance, where the distributions are bounded by censoring upper and lower tails at some mutual cutoff values.
Let $F$ and $G$ be known only via the quantiles at levels $0<\beta_{1}<\beta_{2}<\dots\beta_{k}<1$ with
$F^{-1}(\beta_{i})=F_{i},\qquad G^{-1}(\beta_{i})=G_{i}$
for $i=1,\dots,k$. In what follows, we assume that the distributions have a common bounded support $[l,u]$, and we set $F_{0}=G_{0}=l$, $F_{k+1}=G_{k+1}=u$, $\beta_{0}=0$, and $\beta_{k+1}=1$. Furthermore, we assume that the levels are symmetric about the median level, i.e., $\beta_{i}=1-\beta_{k+1-i}$.
(a) AVM upper bound
(b) AVM lower bound
(c) AVM Shift upper bound
(d) AVM Dispersion upper bound
[…]
The distance between the two distributions is maximized by moving the quantiles in between the known values as far away from one another as possible without violating the monotonicity of a quantile function. Thus, we obtain the distributions $F_{\max},G_{\max}$ with quantile functions [^6] Note that these quantile functions (and the following) are not necessarily left-continuous, as left-continuity may be violated when setting $F_{\max}^{-1}(\beta_{i})=F_{i}$ for example. However, there are left-continuous quantile functions with the known quantiles that are arbitrarily close to these quantile functions. given by
$F_{\max}^{-1}(\alpha)=\begin{cases}F_{i+1}&\text{if }F_{i+1}-G_{i}\geq G_{i+1}-F_{i},\\ F_{i}&\text{otherwise,}\end{cases}\qquad G_{\max}^{-1}(\alpha)=\begin{cases}G_{i}&\text{if }F_{i+1}-G_{i}\geq G_{i+1}-F_{i},\\ G_{i+1}&\text{otherwise,}\end{cases}$
for $\alpha\in(\beta_{i},\beta_{i+1})$ that maximize the distance between distributions subject to having the known quantiles.
Consequently, the AVM between the (unknown) distributions $F$ and $G$ is bounded by the distance between $F_{\max}$ and $G_{\max}$, as illustrated in Figure S5(a). With $\overline{b}_{i}=\beta_{i+1}-\beta_{i}$, we obtain the tight upper bound
$\operatorname{AVM}(F,G)\leq\operatorname{AVM}(F_{\max},G_{\max})=\sum_{i=0}^{k}\overline{b}_{i}\max\{F_{i+1}-G_{i},G_{i+1}-F_{i}\}.$ (S1)
Analogously, we can move the qantiles as close as possible together to obtain a lower bound for the distance. We obtain the distributions $F_{\min},G_{\min}$ with quantiles at levels $\alpha\in(\beta_{i},\beta_{i+1})$ given by
$F_{\min}^{-1}(\alpha)=\begin{cases}F_{i}&\text{if }F_{i}\geq G_{i},\\ \min\{F_{i+1},G_{i}\}&\text{otherwise},\end{cases}\qquad G_{\min}^{-1}(\alpha)=\begin{cases}\min\{G_{i+1},F_{i}\}&\text{if }F_{i}\geq G_{i},\\ G_{i}&\text{otherwise},\end{cases}$
which minimize the distance subject to attaining the known quantiles. The AVM between $F$ and $G$ is bounded by the distance between $F_{\min}$ and $G_{\min}$ from below, as illustrated in Figure S5(b). Thus, we obtain the tight lower bound
$\operatorname{AVM}(F,G)\geq\operatorname{AVM}(F_{\min},G_{\min})=\sum_{i=0}^{k}\overline{b}_{i}\left(F_{i}-\min\{G_{i+1},F_{i}\}+G_{i}-\min\{F_{i+1},G_{i}\}\right).$ (S2)
In a similar spirit, we may define distributions that maximize or minimize a particular component. The plus shift component, $\operatorname{Shift}_{+}^{\operatorname{D}}$, may be bounded (from above) by shifting the quantiles of $F$ up as much as possible and the quantiles of $G$ down as much as possible, as illustrated in Figure S5(c). On the other hand, it may be bounded from below by shifting the quantiles of $F$ down as much as possible and the quantiles of $G$ up as much as possible. Naturally, these distributions also yield bounds for the minus component. Thus, we define the distributions $F_{\uparrow},F_{\downarrow},G_{\uparrow},G_{\downarrow}$ with quantiles at levels $\alpha\in(\beta_{i},\beta_{i+1}),i=0,\dots,k$ given by
$F_{\uparrow}^{-1}(\alpha)=F_{i+1},\quad F_{\downarrow}^{-1}(\alpha)=F_{i},\quad G_{\uparrow}^{-1}(\alpha)=G_{i+1},\quad G_{\downarrow}^{-1}(\alpha)=G_{i},$
respectively. With $\overline{a}_{i}=\overline{b}_{i}$ for $i\neq\frac{k}{2}$ and $\overline{a}_{k/2}=\frac{\overline{b}_{k/2}}{2}$ if $k$ is even, we obtain the bounds for the plus shift component as
$\operatorname{Shift}_{+}^{\operatorname{AVM}}(F,G)$ $\geq\operatorname{Shift}_{+}^{\operatorname{AVM}}(F_{\downarrow},G_{\uparrow})=2\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\overline{a}_{i}\left[\min\{F_{k-i}-G_{k-i+1},F_{i}-G_{i+1}\}\right]_{+},$ (S3)
$\operatorname{Shift}_{+}^{\operatorname{AVM}}(F,G)$ $\leq\operatorname{Shift}_{+}^{\operatorname{AVM}}(F_{\uparrow},G_{\downarrow})=2\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\overline{a}_{i}\left[\min\{F_{k-i+1}-G_{k-i},F_{i+1}-G_{i}\}\right]_{+}.$
The minus component is bounded via the usual symmetry.
The dispersion components are bounded by comparing a distribution with central intervals as wide as possible with a distribution with intervals as narrow as possible, as illustrated in Figure S5(d). We define the distributions $F_{\longleftrightarrow},F_{\rightarrow\leftarrow},G_{\longleftrightarrow},G_{\rightarrow\leftarrow}$ with quantiles at levels $\alpha\in(\beta_{i},\beta_{i+1}),i=0,\dots,k$ given by
$F_{\longleftrightarrow}^{-1}(\alpha)$ $=\begin{cases}F_{i+1},&\alpha>\frac{1}{2},\\ F_{i},&\text{otherwise},\end{cases}\quad F_{\rightarrow\leftarrow}^{-1}(\alpha)=\begin{cases}F_{i},&\alpha>\frac{1}{2},\\ \min\{F_{i+1},F_{n-i}\},&\text{otherwise},\end{cases}$
$G_{\longleftrightarrow}^{-1}(\alpha)$ $=\begin{cases}G_{i+1},&\alpha>\frac{1}{2},\\ G_{i},&\text{otherwise},\end{cases}\quad G_{\rightarrow\leftarrow}^{-1}(\alpha)=\begin{cases}G_{i},&\alpha>\frac{1}{2},\\ \min\{G_{i+1},G_{n-i}\},&\text{otherwise}.\end{cases}$
Thus, we obtain the bounds for the dispersion component as
$\operatorname{Disp}_{+}^{\operatorname{AVM}}(F,G)$ $\geq\operatorname{Disp}_{+}^{\operatorname{AVM}}(F_{\rightarrow\leftarrow},G_{\longleftrightarrow})=\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\overline{a}_{i}\left[(F_{k-i}-G_{k-i+1})-(\min\{F_{i+1},F_{k-i}\}-G_{i})\right]_{+},$ (S4)
$\operatorname{Disp}_{+}^{\operatorname{AVM}}(F,G)$ $\leq\operatorname{Disp}_{+}^{\operatorname{AVM}}(F_{\longleftrightarrow},G_{\rightarrow\leftarrow})=\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\overline{a}_{i}\left[(F_{k-i+1}-G_{k-i})-(F_{i}-\min\{G_{i+1},G_{k-i}\})\right]_{+}.$
Lower and upper bounds for the other distances and their components are obtained analogously, e.g., for the CD, we obtain the bounds
$\operatorname{CD}(F,G)\leq\operatorname{CD}(F_{\max},G_{\max})$
$\hskip 14.22636pt=\sum_{i=j}\overline{b}_{i}\overline{b}_{j}\max\{F_{j+1}-G_{i},G_{i+1}-F_{j}\}+\sum_{i>j}2\overline{b}_{i}\overline{b_{j}}[F_{j}-G_{i+1}]_{+}+\sum_{i<j}2\overline{b}_{i}\overline{b_{j}}[G_{i}-F_{j+1}]_{+},$
$\operatorname{CD}(F,G)\geq\operatorname{CD}(F_{\min},G_{\min})$
$\hskip 14.22636pt=\sum_{i=j}\overline{b}_{i}\overline{b}_{j}\left(F_{j}-\min\{G_{i+1},F_{j}\}+G_{i}-\min\{F_{j+1},G_{i}\}\right)+\sum_{i>j}2\overline{b}_{i}\overline{b_{j}}[F_{j+1}-G_{i}]_{+}+\sum_{i<j}2\overline{b}_{i}\overline{b_{j}}[G_{i+1}-F_{j}]_{+}.$
S4.5 Approximation error
An approximation formula chooses distributions $F_{\approx},G_{\approx}$ that respect the known quantiles, thus producing approximations in between the above bounds. Thus, the approximation error is bounded by the deviation between the lower and upper bounds for the distance or one of its components. For the AVM, we obtain (via a simple case distinction)
$|\operatorname{AVM}(F,G)-\operatorname{AVM}(F_{\approx},G_{\approx})|\leq\operatorname{AVM}(F_{\max},G_{\max})-\operatorname{AVM}(F_{\min},G_{\min})$
$=\sum_{i=0}^{k}\overline{b}_{i}\left(\max\{F_{i+1}-G_{i},G_{i+1}-F_{i}\}-F_{i}+\min\{G_{i+1},F_{i}\}-G_{i}+\min\{F_{i+1},G_{i}\}\right)$
$=\sum_{i=0}^{k}\overline{b}_{i}\left(\max\{F_{i+1}-F_{i}+\min\{F_{i},G_{i+1}\}-G_{i},G_{i+1}-G_{i}+\min\{G_{i},F_{i+1}\}-F_{i}\}\right)$
$\leq\sum_{i=0}^{k}\overline{b}_{i}\left(\max\{F_{i+1}-F_{i}+G_{i+1}-G_{i},G_{i+1}-G_{i}+F_{i+1}-F_{i}\}\right)$
$=\sum_{i=0}^{k}\overline{b}_{i}\left(F_{i+1}-F_{i}+G_{i+1}-G_{i}\right)\leq 2(u-l)\max_{i}\overline{b}_{i}=\mathcal{O}(k^{-1})$
if $\max_{i}\overline{b}_{i}=\mathcal{O}(k^{-1})$ (e.g., for equidistant quantile levels $a_{i}=\frac{1}{k}$).
Similar considerations can be applied to the components. Let $I=\{i\in\{0,1,\dots,\lfloor\tfrac{k}{2}\rfloor\}$ and define
$I_{1}=\{i\in I\mid F_{k-i}-G_{k-i+1}\leq F_{i}-G_{i+1}\},\quad I_{2}=I\setminus I_{1}.$
For the plus shift component of the AVM, we obtain
$|\operatorname{Shift}_{+}^{\operatorname{AVM}}(F,G)-\operatorname{Shift}_{+}^{\operatorname{AVM}}(F_{\approx},G_{\approx})|\leq\operatorname{Shift}_{+}^{\operatorname{AVM}}(F_{\uparrow},G_{\downarrow})-\operatorname{Shift}_{+}^{\operatorname{AVM}}(F_{\downarrow},G_{\uparrow})$
$=2\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\overline{a}_{i}\left(\left[\min\{F_{k-i+1}-G_{k-i},F_{i+1}-G_{i}\}\right]_{+}-\left[\min\{F_{k-i}-G_{k-i+1},F_{i}-G_{i+1}\}\right]_{+}\right)$
$\leq 2\sum_{i\in I_{1}}\overline{a}_{i}\left(\left[F_{k-i+1}-G_{k-i}\right]_{+}-\left[F_{k-i}-G_{k-i+1}\right]_{+}\right)+2\sum_{i\in I_{2}}\overline{a}_{i}\left(\left[F_{i+1}-G_{i}\right]_{+}-\left[F_{i}-G_{i+1}\right]_{+}\right)$
$\leq 2\sum_{i\in I_{1}}\overline{a}_{i}\left(F_{k-i+1}-F_{k-i}+G_{k-i+1}-G_{k-i}\right)+2\sum_{i\in I_{2}}\overline{a}_{i}\left(F_{i+1}-F_{i}+G_{i+1}-G_{i}\right)$
$\leq 2\sum_{i=0}^{k}\overline{b}_{i}\left(F_{i+1}-F_{i}+G_{i+1}-G_{i}\right)\leq 4(u-l)\max_{i}\overline{b}_{i}=\mathcal{O}(k^{-1})$
if $\max_{i}\overline{b}_{i}=\mathcal{O}(k^{-1})$.
For the dispersion component of the AVM, we obtain
$|\operatorname{Disp}_{+}^{\operatorname{AVM}}(F,G)-\operatorname{Disp}_{+}^{\operatorname{AVM}}(F_{\approx},G_{\approx})|\leq\operatorname{Disp}_{+}^{\operatorname{AVM}}(F_{\longleftrightarrow},G_{\rightarrow\leftarrow})-\operatorname{Disp}_{+}^{\operatorname{AVM}}(F_{\rightarrow\leftarrow},G_{\longleftrightarrow})$
$=\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\overline{a}_{i}\left(\left[(F_{k-i+1}-G_{k-i})-(F_{i}-\min\{G_{i+1},G_{k-i}\})\right]_{+}-\left[(F_{k-i}-G_{k-i+1})-(\min\{F_{i+1},F_{k-i}\}-G_{i})\right]_{+}\right)$
$\leq\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\overline{a}_{i}\left((F_{k-i+1}-G_{k-i})-(F_{i}-\min\{G_{i+1},G_{k-i}\})-(F_{k-i}-G_{k-i+1})-(\min\{F_{i+1},F_{k-i}\}-G_{i})\right)$
$=\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\overline{a}_{i}\left(F_{k-i+1}-F_{k-i}+\min\{F_{i+1},F_{k-i}\}-F_{i}+G_{k-i+1}-G_{k-i}+\min\{G_{i+1},G_{k-i}\}-G_{i}\right)$
$=\sum_{i=0}^{k}\overline{b}_{i}(F_{i+1}-F_{i}+G_{i+1}-G_{i})\leq 2(u-l)\max_{i}\overline{b}_{i}=\mathcal{O}(k^{-1})$
if $\max_{i}\overline{b}_{i}=\mathcal{O}(k^{-1})$.
Analogous considerations apply to the approximation error for the other distances and their components. For instance for the CD, we obtain
$|\operatorname{CD}(F,G)-\operatorname{CD}(F_{\approx},G_{\approx})|\leq\operatorname{CD}(F_{\max},G_{\max})-\operatorname{CD}(F_{\min},G_{\min})$
$\leq\sum_{i=j}\overline{b}_{i}\overline{b}_{j}(F_{j+1}-F_{j}+G_{i+1}-G_{i})+\sum_{i\neq j}2\overline{b}_{i}\overline{b_{j}}(F_{j+1}-F_{j}+G_{i+1}-G_{i})$
$\leq 4k(u-l)(\max_{i}\overline{b_{i}})^{2}=\mathcal{O}(k^{-1})$
as long as $\max_{i}\overline{b}_{i}=\mathcal{O}(k^{-1})$.
@@
For real numbers $A,B\in\mathbb{R}$ and $A]_{+}:=\max(A,0)$, we have
@@
  • If $A,B>0$, the statement is immediate as $A]_{+}+[B]_{+}=A+B$ and $A+B]_{+}+[\min\{A,-B\}]_{+}+[\min\{-A,B\}]_{+}=A+B+0+0$.
  • If $A,B\leq 0$, then $A]_{+}+[B]_{+}=0+0$ and $A+B]_{+}+[\min\{A,-B\}]_{+}+[\min\{-A,B\}]_{+}=0+0+0$.
  • If $A>0$ and $B\leq 0$, we get that $A]_{+}+[B]_{+}=A+0=A$. If additionally $A\geq-B$, we get $A+B]_{+}+[\min\{A,-B\}]_{+}+[\min\{-A,B\}]_{+}=(A+B)+(-B)+0=A$. Similarly, if A<BA<-B, we get $A+B]_{+}+[\min\{A,-B\}]_{+}+[\min\{-A,B\}]_{+}=0+A+0=A$.
@@
where the third equality is obtained by a simple change of variables and the last equality is obtained by applying Lemma S5.1 to $A(\alpha)]_{+}+[-B(\alpha)]_{+}$ and $A(\alpha)]_{+}+[B(\alpha)]_{+}$.
@@
[…]
@@
[…]
@@
[…]
@@
Notice that for the two minus components with subscript ‘-’, changing the roles of $F$ and $G$ merely changes the sign of $A$ and $B$. ∎
@@
We start by proving Propositions 3.8 and 3.9 as these help to simplify the proofs of some of the basic properties given in Section 3.1. Notice that the proofs of Propositions 3.8 and 3.9 do not require any of the results from Section 3.1 apart from the obvious symmetry given by Proposition 3.1.
@@
    From Proposition 3.9 (shown above), it follows that $\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)=0$ is zero as well for all $p\in[1,\infty)$.
@@
By Propositions 3.8 and 3.9 (proved above), it suffices to show the equivalences in (a) and (b) for the AVM.
@@
    is zero. Proposition 3.9 implies that $\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)=0$ for all $p\in[1,\infty)$.
@@
[…]
@@
[…]
@@
To prove Theorem 3.11, we need the following lemma.
Lemma S6.1.
Let $q>p\geq 1$. Then, the inequality
$a^{p}+(a-2)^{p}\leq a^{q}+(a-2)^{q}$ (S8)
holds for $a\geq 2$.
Proof.
The function $g(p)=a^{p}+(a-2)^{p}$ has first and second derivative
$g^{\prime}(p)=a^{p}\ln(a)+(a-2)^{p}\ln(a-2),\quad\text{and}\quad g^{\prime\prime}(p)=a^{p}\ln(a)^{2}+(a-2)^{p}\ln(a-2)^{2},$
respectively. For the inequality to hold, it suffices to show that the first derivative $g^{\prime}$ is nonnegative (for $p\geq 1$) and thus the function $g$ is increasing. As the second derivative is clearly positive, the function is (strictly) convex. Thus, it suffices to show that $g^{\prime}(1)=a\ln(a)+(a-2)\ln(a-2)=h(a)\geq 0$ for $a\geq 2$. The function $h$ has first and second derivative
$h^{\prime}(a)=\ln(a)+\ln(a-2)+2\quad\text{and}\quad h^{\prime\prime}(a)=\frac{1}{a}+\frac{1}{a-2},$
respectively. It is easy to see that $h$ is convex on $2,\infty)$ with a minimum at $a_{0}=1+\sqrt{1+e^{-2}}$. Therefore, we have $g^{\prime}(1)=h(a)\geq h(a_{0})>0$ as $h(a_{0})\approx 1.32$. ∎
@@
Let $\widetilde{m}=m_{F}-m_{G}$ be the difference between the central medians of $F$ and $G$ (as defined prior to Proposition 3.3). If $\widetilde{m}=0$, then both shift components are zero by Proposition 3.3 and inequality (23) is trivial. We assume w.l.o.g. that $\widetilde{m}>0$ (otherwise, we exchange $F$ and $G$). Then, the minus shift components are zero by Proposition 3.3. Furthermore, we define $f(\tau)=F^{-1}(\tau)-G^{-1}(\tau)$ for $\tau\in(0,1)$. By symmetry of $F$ and $G$, the function $f$ is (almost surely) symmetric as well, that is,
$f(\tau)=2\widetilde{m}-f(1-\tau)$ (S9)
holds for almost all $\tau\in(0,1)$. Let $A=\{\tau\mid f(1-\tau)>f(\tau)\geq 0\}\cup\{\tau<\tfrac{1}{2}\mid f(1-\tau)=f(\tau)\geq 0\}$. Then, the plus shift component of the $p$-Wasserstein distance is twice the integral of $f$ over $A$ as a change of variables yields $\operatorname{Shift}_{+}^{\operatorname{WD}_{p}}(F,G)=\int_{0}^{1}\left[\min\left\{f(\tau)^{[p]},f(1-\tau)^{[p]}\right\}\right]_{+}\,\mathrm{d}\tau=2\int_{A}|f(\tau)|^{p}\,\mathrm{d}\tau$. Let $\overline{A}=\{1-\tau\mid\tau\in A\}$, $B=\{\tau\mid f(\tau)\geq 0>f(1-\tau)\}$, and $\overline{B}=\{1-\tau\mid\tau\in B\}$. By the (almost sure) symmetry of $f$, the values $\tau$ such that $0>f(\tau)$ and $0>f(1-\tau)=2\widetilde{m}-f(\tau)$ form a null set in $(0,1)$, and so does the complement of the disjoint union $A\cup\overline{A}\cup B\cup\overline{B}$ in $(0,1)$. Note that the symmetry (S9) of $f$ yields the inequalities
$\widetilde{m}$ $\geq f(\tau)$ $\quad\text{for almost all }\tau\in A$ (S10)
$f(1-\tau)$ $\geq\widetilde{m}$ $\quad\text{for almost all }\tau\in A$ (S11)
$f(\tau)$ $\geq 2\widetilde{m}$ $\quad\text{for almost all }\tau\in B$ (S12)
We use these a.s. (almost sure with respect to the Lebesgue measure) inequalities to derive the inequality